OpenStudy (anonymous):

Researchers at the University of British Columbia have measured muscle reaction times of 0.1 seconds for Olympic sprinters. Suppose the human chain consisted entirely of Olympic sprinters with an average arm span of 6 feet. It would take the signal ____ days to travel all the way around the world. Round your answer to the nearest tenth of a day and enter only the number. (Hint: First compute the number of people required to circle the globe, using 24,000 miles for the circumference of the Earth and 5280 feet per mile.)

OpenStudy (mertsj):

\[24000 mi \times\frac{5280 ft}{1 mi} \times\frac{1 person}{6 ft} \times \frac{.1 \sec}{1person}\times \frac{1 hour}{3600 \sec} \times\frac{1 day}{24 hours}= 244.4days\]

OpenStudy (anonymous):

Thank you!